OpenStudy (anonymous):

A hypothetical planet has a mass 1.90 times that of Earth, but the same radius. What is g near its surface?

OpenStudy (anonymous):

Use that the force is \[F=G*m*m p/r ^{2}\] with G=6.67e-11, m the mass of your point and mp the mass of the planet. Then consider you want your force to be some \[F=m*g\]. So : \[g = G*m p/r ^{2}\]

OpenStudy (anonymous):

When I plug in my values I get: \[6.673*10^-11Nm^2/kg^2*1.136kg/(6.38*10^3km)^2=1.86m/s^2\] I must have gotten the wrong mass of the planet because its saying my answer is wrong. I did this to get the mass of the planet: \[5.98*10^{24}*1.90=1.136kg\]

OpenStudy (anonymous):

Is this the answer: \[g=Gm _{p}/r^2(6.61*10^-11Nm^2/kg^2)(5.98*10^24kg)(1.1362*10^25)/(6.38*10^3km)^2\]=\[18.62g\]

OpenStudy (anonymous):

nevermind the answer was 18.6m/s